# Gamma-ray Bursts as Distance Indicators by a Machine Learning Approach

### v1

@@ A.3.1 Errorbars on predicted redshift and the catastrophic outliers
where $*\sigma*$ is the standard deviation of the residual Gaussian fit obtained in the $\log(z+1)$ scale.
@@ A.4 Optimal Transport bias correction
When the predictions of an ML model are skewed towards a particular range of values, then such a model is said to be biased. Bias correction techniques attempt to minimize such biases in the model predictions. Bias is defined as the mean of the difference between the model’s prediction and the actual values. Biased results from ML models generally arise due to imbalanced training data, where more observations in a particular range are present compared to others. In the case of GRBs, there are more observations at low redshifts compared to high redshift ones. Thus, an ML model trained on such data will be more inclined toward underestimating the redshift of GRBs.
@@
It should be noted here that the bias correction is applied following the removal of the catastrophic outliers, specifically the points outside of the blue cone in the left panel of Fig. 9.
The right panel shows the prediction after removing the catastrophic outliers and the bias correction. We also stress that if we remove the catastrophic outliers additionally after the bias correction, we obtain an even higher correlation coefficient between the $z_{pred}$ and $z_{obs}$ 0.94. The bias correction can also be applied in the generalization sample as long as the training set and the generalization sample are drawn by the same parent population.
@@ A.5 EP method
[…]
@@
Figure 10: Left Panel: Luminosity vs. redshift for 103 GRBs. The luminosities calculated using $z_{obs}$ and $z_{pred}$ are shown with blue and red points, respectively. The orange and purple points represent the luminosities of $z_{obs}$ and $z_{pred}$, respectively that are lower than the luminosity threshold given by the flux limit of $10^{-13}erg/cm^{2}$. Right panel: $\tau$ versus $k$, where $k$ defines the power of the $g(z)$ function for the luminosity evolution. The red curve shows results for the GRBs with observed $z_{obs}$, while the blue curve yields the corresponding function using $z_{pred}$. The $\tau=0$ values are obtained at $k=3.07$ for the GRBs with $z_{obs}$ and $3.06$ for $z_{pred}$.

### v2

@@ A.3.1 Errorbars on predicted redshift and the catastrophic outliers
where $\sigma*$ is the standard deviation of the residual Gaussian fit obtained in the $\log(z+1)$ scale.
@@ A.4 Optimal Transport bias correction
When the predictions of a statistical learning model are skewed towards a particular range of values, then such a model is said to be biased. Bias correction techniques attempt to minimize such biases in the model predictions. Bias is defined as the mean of the difference between the model’s prediction and the actual values. Biased results from statistical learning models generally arise due to imbalanced training data, where more observations in a particular range are present compared to others. In the case of GRBs, there are more observations at low redshifts compared to high redshift ones. Thus, a statistical learning model trained on such data will be more inclined toward underestimating the redshift of GRBs.
@@
[…]
The right panel shows the prediction after removing the catastrophic outliers and the bias correction. We also stress that if we remove the catastrophic outliers additionally after the bias correction, we obtain an even higher correlation coefficient between the $z_{pred}$ and $z_{obs}$ of 0.94. The bias correction can also be applied in the generalization sample as long as the training set and the generalization sample are drawn by the same parent population.
@@ A.5 EP method
The EP method requires a limiting flux, $F_{lim}$. For $F_{lim}$ we chose $L_{lim}=4\pi D_{L}^{2}(z)\,F_{lim}K$, where $D_{L}$ is the luminosity distance and $K$ is the K-correction for cosmic expansion (Bloom et al., 2001), which gives the minimum observed luminosity for a given redshift. In the left panel of Fig. 10, we show the limiting luminosity for $K=1$, not to have fuzzy boundaries, but for an appropriate evaluation of the luminosity evolution, we assign to each GRB its own K correction. We have investigated several limiting fluxes to determine a good representative value while keeping an adequate sample size. We have finally chosen the limiting flux $F_{lim}=1.25\times 10^{-14}erg/cm^{-2}$, for both the observed and predicted redshift, which allows $97$ GRBs. Limiting luminosity and luminosity of GRBs from the observed (red circles) and predicted redshifts (blue circles) are shown on the left panel of Fig. 10. The orange and purple circles show the removed GRBs based on the chosen flux limit.
@@
Figure 10: Left Panel: Luminosity vs. redshift for 103 GRBs. The luminosities calculated using $z_{obs}$ and $z_{pred}$ are shown with blue and red points, respectively. The orange and purple points represent the luminosities of $z_{obs}$ and $z_{pred}$, respectively that are lower than the luminosity threshold given by the flux limit of $1.25\times 10^{-14}erg/cm^{2}$. The arrows connect a GRB’s $z_{obs}$ and luminosity value with its predicted values. Right panel: $\tau$ versus $k$, where $k$ defines the power of the $g(z)$ function for the luminosity evolution. The red curve shows results for the GRBs with observed $z_{obs}$, while the blue curve yields the corresponding function using $z_{pred}$. The $\tau=0$ values are obtained at $k=3.07$ for the GRBs with $z_{obs}$ and $3.06$ for $z_{pred}$.